PDS_VERSION = PDS3 RECORD_TYPE = STREAM OBJECT = TEXT PUBLICATION_DATE = 1997-09-11 NOTE = "MGS RST Instrument Health Report HEA7254A.TXT" END_OBJECT = TEXT END Miscellaneous ============= From 5 PM Tuesday until 9:15 AM this morning we were unable to connect to the Query Server. A call to Ben Jai this morning seems to have fixed the problem. There are no spacecraft or DSN Monitor data for the past two days; whether we will be able to recover these depends on the volume of data and the number of competing activities over the next few days. There was no report yesterday because of the MGS RST meeting at JPL. Engineering Data ================ No new data. Open Loop Data ============== Data Collection --------------- About an hour of open loop data were obtained during pressurization on Tuesday. The DSP ran longer, but observing two-way transmissions. Details are below for the segment transferred to Stanford: YYYY/DDD DSS QUERY SAMP BW DSP BIT TRK REC RECS RECS MAX START STOP RATE MODE RESN MODE LEN MSSG SNR -------- --- -------- -------- ---- ---- ---- ---- ---- ---- ----- ---- ----- 1997/252 65 14:00:00 15:12:11 5000 2500 2 12 1-W 1666 21654 2 33. Anomalies --------- There were a number of attenuator changes, as operators tracked changes in spacecraft configuration (HGA vs LGT, for example). There appear to be some quantization errors in the sample distribution (effectively 11-bit samples as opposed to the 12-bits that are taken). Signal levels were low; no spurs are visible in quick-look power spectra. Overall, the anomalies in these data were minor. Analysis -------- Joe Twicken obtained X-band USO frequency 8423152966.841 Hz from the data; this falls precisely on the linear drift of +0.182 Hz/day observed since mid- May. Closed Loop Data ================ Data Collection --------------- We have obtained the following files from OSCAR and EPIC: Original File Name Name on SOPC ------------------------------ ------------ ATDF: ODF: 97252C253.DAT;1 7252253A.ODF 97253C254.DAT;1 7253254A.ODF 97254C254.DAT;1 7254254A.ODF IONCAL: TROPCAL: WEATHER: WEATHER_10_LATEST.TXT;71 70012531.WEA WEATHER_40_LATEST.TXT;70 70012534.WEA WEATHER_60_LATEST.TXT;70 70012536.WEA EOP: 970909-eop.long 20207287.EOP 970909-eop.short 62877287.EOP Ancillary Data ============== Data Collection --------------- No new files